A loop of wire in the form of a square 2.20 m on each side, its plane makes an angle of 36.0° with a uniform magnetic field of 0.80 T. What is the magnetic flux through the loop?

( I give thumbs up if correct right away, so rlly appreciate ur help :)
Answer #1

Magnetic flux = NIAB Sin(theta)

N=1, I= current should be given, A= 2.2^2 =4.84 m^2, B= 0.8 T,

Theta = 36 degrees

Flux = 2.276 I (I = current)